OpenStudy (anonymous):

evaluate the formula A=pie radius2 for pie 22/7 and r=7/4

OpenStudy (anonymous):

i think you just have to substitute the values into the formula

OpenStudy (carson889):

First off, the number 3.14 is spelt pi not pie, as pie is a food and pi is most certainly not edible. If you're evaluating for r = 7/4, first you square it so r^2 = 49/16, now just multiply by pi so it equals approximately 9.62. Now for r = 22/7, r^2 = 484/49. Now multiply that number by pi to get approximately 31.0. Sorry didn't mean to be knit picky in the beginning :P

OpenStudy (anonymous):

that is ok i meant to pi just accidentally put the e on the end of it